OpenStudy (anonymous):

x^2y+x+sin^2y=4 how can I find the derivative implicitly?

OpenStudy (phi):

treat both x and y as variables. if you take the derivative with respect to x you have \[ \frac{d}{dx} x = \frac{dx}{dx}= 1 \\ \text{ and }\\ \frac{d}{dx} y = \frac{dy}{dx} \] use those ideas and the usual "rules": chain rule, derivative of a product, etc.